OpenStudy (anonymous):

Factor the trinomial... V^2-4vy-45y^2

OpenStudy (anonymous):

** Hint -9*5=-45 -9+5=4

OpenStudy (anonymous):

now what do i do with that put it into the problem?

OpenStudy (bimbim2697):

-9+5=4

OpenStudy (anonymous):

so that is the answer... it is not 2 different trinomials?